Not sure if I'm figuring this one right.
y=mx-5, for m

Answer by tutorcecilia(2152)

y=mx-5
y+5=mx-5+5
y+5=mx
(y+5)/x=mx/x
(y+5)/x=m
.
Check by plugging (y+x)=m back into the original equation:
y=mx-5
y=(y+5/x)(x) -5 [Simpllify]
y=y+5-5
y=y [checks out]